In watching a transverse wave go by, an observer notes that 13 crests go by in a time of 3.0 s. If the distance between two successive crest is measured to be 0.75 m, (a) what is the speed of the wave? (b) What is its period?


1
Expert's answer
2022-05-02T13:43:43-0400

"T=\frac tn=0.23~s,"

"v=\frac lT=3.25~\frac ms."